OpenStudy (anonymous):

Will give Medal You have a $40 coupon from the manufacturer good for the purchase of a cell phone. The store where you are purchasing your cell phone is offering a 30% discount on all cell phones. Let x represent the regular price of the cell phone.

OpenStudy (anonymous):

a) Suppose only the 30% discount applies. Find a function f that models the purchase price of the cell phone as a function of the regular price x.

OpenStudy (studygurl14):

\[f(x) = x - (x \times .3)\]

OpenStudy (anonymous):

@StudyGurl14 (b) Suppose only the $40 coupon applies. Find a function g that models the purchase price of the cell phone as a function of the sticker price x.

OpenStudy (studygurl14):

\[g(x) = x - 40\]

OpenStudy (anonymous):

last one (c) If you can use the coupon and the discount, then the purchase price is either f compose g(x) or g compose f(x), depending on the order in which they are applied to the price. Find both f compose g(x) and g compose f(x).